OpenStudy (anonymous):

Find sin2x cos2x tan2x from the following: csc(x)=4, tan(x) < 0

OpenStudy (anonymous):

is the answer cos2x = 3/4 sin2x = rad(7)/4 tan2x = rad(7)/3 ?

OpenStudy (anonymous):